OpenStudy (anonymous):

help!!! Find an equation for the nth term of the arithmetic sequence. -3, -5, -7, -9, ...

OpenStudy (anonymous):

Still struggling I see :)

OpenStudy (anonymous):

You use the same formula as for the other one. Un = a + d(n-1)

OpenStudy (anonymous):

yes I keep trying to use a formula and my answer doesn't match up with the choices

OpenStudy (anonymous):

This one is quite straight forward. The first term is -3 and the common difference is -2.

OpenStudy (anonymous):

I dont understand what I plug in for n, d, a,…..

OpenStudy (anonymous):

N is the term you want to find the value of. Say you want to find the value of the 27th term, you would plug 27 in for every "n" in the formula.

OpenStudy (anonymous):

a is the first term of the sequence. In this case it is -3.

OpenStudy (anonymous):

d is the common difference between each terms. This can be found by subtracting the first term from the second. In this case -5 -3 = -2.

OpenStudy (anonymous):

So you plug these values into the formula making it Un = -3 + -2(n-1)

OpenStudy (anonymous):

If you want to find the 27th term as I mentioned earlier, it would be U27 = -3 + -2(27-1)

OpenStudy (anonymous):

I understand what you are doing now, it is much clearer but why 27?

OpenStudy (anonymous):

Just a random number to show how it works, I could have picked any.

OpenStudy (jdoe0001):

\(\begin{array}{cccccclll} a_1\\ \downarrow \\ -3& -5& -7& -9& ...\\ \hline\\ &-3-2&(-3-2)-2&(-3-2-2)&-2\\ &&&&\uparrow \\ &&&&d \end{array}\quad \Large a_n=a_1+(n-1)d\)

OpenStudy (anonymous):

ohh okay I got it thanks you!

OpenStudy (anonymous):

wait!! what if all of the numbers in the sequence were positive? Would i use n+1 instead of n-1??

OpenStudy (jdoe0001):

then "d" will be positive, (n-1) stays always the same

OpenStudy (anonymous):

thanks!!

OpenStudy (jdoe0001):

yw